# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

### v1

@@ Abstract
Multivariable time series forecasting methods can integrate information from exogenous variables, leading to significant prediction accuracy gains. Transformer architecture has been widely applied in various time series forecasting models due to its ability to capture long-range sequential dependencies. However, a naïve application of transformers often struggles to effectively model complex relationships among variables over time. To mitigate against this, we propose a novel architecture, namely the Spectral Operator Neural Network (Sonnet). Sonnet applies learnable wavelet transformations to the input and incorporates spectral analysis using the Koopman operator. Its predictive skill relies on the Multivariable Coherence Attention (MVCA), an operation that leverages spectral coherence to model variable dependencies. Our empirical analysis shows that Sonnet yields the best performance on $34$ out of $47$ forecasting tasks with an average mean absolute error (MAE) reduction of $1.1\%$ against the most competitive baseline (different per task). We further show that MVCA—when put in place of the naïve attention used in various deep learning models—can remedy its deficiencies, reducing MAE by $10.7\%$ on average in the most challenging forecasting tasks.
Source code: github.com/ClaudiaShu/Sonnet
@@ 1 Introduction
Multivariable time series (MTS) forecasting methods learn from multiple input variables to predict a single target variable [17]. MTS models are deployed in many real-life applications, such as financial modelling [35, 3, 46, 54], numerical weather prediction [13, 55, 8], and computational epidemiology [11, 42, 7, 31]. Recent machine learning models for time series forecasting tend to emphasise multivariate formulations [34, 58, 24, 2, 9]. While such models provide important insights for integrating traditional time series forecasting into deep learning techniques—e.g. capturing seasonality [23], frequency-domain methods [62, 63], and autoregression [58, 34]—they often overlook the benefits of leveraging external information through exogenous indicators.
@@
In fact, empirical evidence suggests that capturing dependencies from external variables can increase the risk of overfitting for complex models [34]. Moreover, existing methods that capture inter-variable dependencies [60, 51] have, in certain forecasting tasks, been outperformed by models that do not [23, 26]. However, this can be attributed to the choice of forecasting benchmarks, which, in many occasions, are characterised by strong seasonal patterns [23] (see also Appendix E.4), that can be effectively modelled using simpler autoregressive methods [58, 23]. Additionally, models that do not capture inter-variable dependencies benefit from learning using longer look-back windows without significantly increasing modelling complexity and computational cost [15], making them less likely to overfit. Nevertheless, these empirical outcomes are not necessarily valid for multivariable forecasting scenarios, and especially for tasks where exogenous variables have strong predictive power. In such cases, explicitly modelling interactions across variables facilitates better performance [52, 43].
@@
Prior work has explored various ways in capturing inter-variable dependencies for multivariate or MTS forecasting tasks. Methods applying attention [48] across variables [25, 18] can capture nonlinear dependencies, but disrupt temporal information by embedding sequences along the time dimension. Crossformer [60] attempts to rectify this by proposing a modified transformer structure where input is split into time series patches allowing to capture dependencies across both the time and variable dimensions at the sub-sequence level. ModernTCN [27], on the other hand, uses a convolutional kernel over both the variable and time dimension to jointly capture the inter- and intra-variable dependencies. However, both Crossformer and ModernTCN suffer from GPU computational overhead as the number of exogenous variables or the length of the input time series increases [43, 64]. TimeXer [52] models dependencies between exogenous variables and the target variable separately, and then joins their learned embeddings with a cross-attention module. DeformTime [43] yields superior results by using deformable attention to incorporate information from exogenous variables at different time steps. However, it only captures inter-variable dependencies within the reception field of a convolutional kernel, which limits the exploration of a wider range of exogenous variables.
@@
Frequency-based analysis, including using Fourier [6, 44] or wavelet transform [47, 59, 4, 12, 56], has been widely used in statistical approaches for identifying periodic patterns [37] as well as in machine learning methods for forecasting [20, 62, 63, 36]. These methods compress temporal information (due to the Fourier transform) while some of them focus on capturing intra-variable dependencies [62, 63]. Wavelet transform, albeit dependent on the chosen mother wavelet [10, 32], can maintain both time and frequency information. In the frequency domain, spectral coherence [53] serves as a powerful tool for capturing the correlation between variables at different frequencies [45, 30]. To improve input projections into the frequency domain, Lange et al. [20] used a Koopman operator [29, 40, 5, 24], i.e. a spectral function that enables linear modelling of nonlinear changes to better address nonlinearities. More recently, AdaWaveNet [56] decomposed the input series into seasonal and long-term components, and used the wavelet transform to capture periodic information. This method accounts for inter-variable dependencies only over the seasonal components.
@@
Motivated by the aforementioned remarks, we propose Spectral Operator Neural Network (Sonnet), a model that captures time series dependencies in the spectral domain using a learnable wavelet transform. Sonnet captures both intra- and inter-variable dependencies using a novel frequency-domain Multivariable Coherence Attention (MVCA) layer. Furthermore, it deploys a learnable Koopman operator that facilitates linearised transitions of temporal states [22]. MVCA demonstrates stand-alone effectiveness when integrated into existing architectures, outperforming naïve attention and other modified attention mechanisms. Our contributions can be summarised as follows:
@@
    We assess forecasting accuracy on carefully curated MTS data sets. This includes established benchmarks complemented by tasks on weather forecasting, influenza prevalence, electricity consumption, and energy prices. The weather prediction and influenza prevalence tasks support a more thorough evaluation because they include a substantial amount of exogenous variables, multiple years ($\geq 10$) for training, and multiple test periods ($\geq 3$).
@@
    Sonnet reduces mean absolute error (MAE) by $1.1\%$ on average compared to the most performant baseline (varies each time). In the more challenging tasks of influenza and weather modelling, MAE is reduced by $3.5\%$ and $2\%$, respectively. Performance gains persist as the forecasting horizon increases, demonstrating the effectiveness of Sonnet in longer-term forecasting.
@@ 2 MTS forecasting task definition
We focus exclusively on multivariable time series (MTS) forecasting, whereby multiple input variables are used to predict a single target variable. We note that although for some baseline models multiple output variables may be present (multivariate forecasting), our evaluation is restricted to the prediction of one specific output. All models are trained and evaluated under a rolling window setup, with a look-back window and a forecasting horizon of $L$ and $H$ time steps, respectively. Specifically, at each time step $t$, the $C$ observed exogenous variables over $L$ past time steps, $\{t-L+1,\dots,t-1,t\}$, are captured in an input matrix $\mathbf{X}_{t}\in\mathbb{R}^{L\times C}$. The autoregressive signal for the target (endogenous) variable is denoted by $\mathbf{y}_{t-\delta}\in\mathbb{R}^{L}$; this encompasses time steps $\{t-\delta-L+1,\dots,t-\delta-1,t-\delta\}$, where $\delta\in\mathbb{N}_{0}$ is an optional delay applied when the target variable is observed with a temporal lag. We capture the exogenous and endogenous input variables in $\mathbf{Z}_{t}=[\mathbf{X}_{t},\mathbf{y}_{t-\delta}]\in\mathbb{R}^{L\times(C+1)}$. The goal is to predict the target variable at time step $t+H$, where $H$ denotes the forecasting horizon. Hence, the output of a forecasting model is denoted by $\mathbf{y}_{t+H}\in\mathbb{R}^{H}$ and holds forecasts for time steps $\{t+1,\dots,t+H-1,t+H\}$. For models that conduct multivariate forecasting, the output includes predictions for all covariates and hence is denoted by $\mathbf{Y}_{t+H}=[\mathbf{X}_{t+H},\mathbf{y}_{t+H}]\in\mathbb{R}^{H\times(C+1)}$. The forecasting task is to learn $f:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Performance is measured based on the endogenous forecast at time step $t+H$, i.e. the last (temporally) element of $\mathbf{y}_{t+H}$, $y_{t+H}\in\mathbb{R}$. For notational simplicity, we omit temporal subscripts and use $\mathbf{Z}$ for $\mathbf{Z}_{t}$, $\mathbf{y}$ for $\mathbf{y}_{t-\delta}$, and $\mathbf{X}$ for $\mathbf{X}_{t}$.
@@ 3 Spectral coherence with Sonnet
In this section, we provide a detailed description of our proposed model, Sonnet. It operates in the spectral domain via a learnable wavelet transform and uses a Multivariable Coherence Attention (MVCA) module to capture both inter- and intra-variable dependencies. We further use a Koopman projection layer that enables stable temporal evolution via a learned linear operator.
3.1 Joint embedding of exogenous and endogenous variables
@@ 3.2 Learnable wavelet transform
We then transform the time series embedding into the wavelet space, i.e. a space that contains both time and frequency information. Specifically, after obtaining the input time series embedding $\mathbf{E}$, we first define a set of $K$ learnable wavelet transformations (also referred to as atoms), with the $k$-th atom held in a matrix $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ derived by
@@
where $\mathbf{t}\in\mathbb{R}^{L}$ is a row vector capturing normalised time steps with $\mathbf{t}_{i}=i/(L-1)$ for $i=0,\dots,L-1$, and $\mathbf{w}_{\alpha}$, $\mathbf{w}_{\beta}$, and $\mathbf{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable weight vectors that control the shape of the wavelet, each initialised randomly from a normal distribution. In particular, $\mathbf{w}_{\alpha}$ controls the width of the Gaussian envelope, and $\mathbf{w}_{\beta}$, $\mathbf{w}_{\gamma}$ respectively determine the linear or quadratic frequency modulation of the generated cosine waveforms. This formulation enables the atoms to adapt to localised time-frequency structures in the data. The time series embedding $\mathbf{E}\in\mathbb{R}^{L\times d}$ is then transformed into the wavelet space by projecting it onto each of the wavelet atoms $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ using $\mathbf{P}_{k}=\mathbf{E}\odot\mathbf{M}_{k}^{\top}$, where $\mathbf{P}_{k}\in\mathbb{R}^{L\times d}$ denotes the embedding’s projection for the $k$-th atom. The transformed wavelet across all $K$ atoms is denoted by $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$. The aforementioned steps help to preserve temporal structure while decomposing the input into multi-resolution time-frequency components using adaptive wavelet transforms that can capture both short- and long-term patterns in the data.
@@ 3.3 Multivariable Coherence Attention (MVCA)
Given the obtained input embedding matrix in wavelet space $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$, where $L$ and $d$ are the time and variable dimension, $K$ is the number of wavelets after transformation, we first linearly project it to query, key and value embeddings, denoted as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}\in\mathbb{R}^{K\times L\times d}$, using weight matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ respectively, as in $\mathbf{Q}=\mathbf{P}\mathbf{W}_{q}$. We then consider the embeddings from different wavelet transformations (indexed by $k$) as separate attention heads. Specifically, for each attention head, we obtain the sub-tensor of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, denoted as $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}\in\mathbb{R}^{L\times d}$, as the query, key, and value embeddings. Therefore, each head captures a distinct subspace of the original embedding, where subspaces correspond to different wavelet transformations of the input. Including a multi-head structure enables the model to learn diverse dependencies in parallel [48].
We then apply a Fast Fourier Transform (FFT) along the variable dimension of each transformer head to transfer the query and key embeddings to the frequency domain, i.e. $\mathbf{Q}_{f}=\texttt{FFT}\left(\mathbf{Q}_{h}\right)$ and $\mathbf{K}_{f}=\texttt{FFT}\left(\mathbf{K}_{h}\right)$, with both $\mathbf{Q}_{f}$ and $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$, where $\ell=\lfloor\frac{d}{2}\rfloor+1$. Each frequency bin in the transformed data ($\mathbf{Q}_{f}$, $\mathbf{K}_{f}$) contains information from all original inputs to the FFT [21], capturing both fine-grained (high-frequency) and global (low-frequency) patterns. The cross-spectral density $\mathbf{P}_{qk}\,\in\,\mathbb{C}^{L\times\ell}$ and power-spectral densities $\mathbf{P}_{qq}$, $\mathbf{P}_{kk}\,\in\,\mathbb{R}^{L\times\ell}$ are obtained as follows:
$\mathbf{P}_{qk}=\mathbf{Q}_{f}\odot\mathbf{K}_{f}^{\ast}\,\,,\quad\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}\,\,,\quad\text{and}\,\quad\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}\,\,,$ (2)
@@
where $\epsilon=10^{-6}$ mitigates division by 00. $\mathbf{C}_{qk}$ captures the linear dependency between sequences across frequency bands. Therefore, the coherence here assigns higher importance to the time steps where the query and key hold more similar averaged values across multiple frequencies.
Following the common design of attention layers [48], we scale, normalise, and regularise $\mathbf{C}_{qk}$, i.e. $\mathbf{A}_{h}=\texttt{Dropout}(\texttt{Softmax}(\mathbf{C}_{qk}/\sqrt{d}))$. The attention weights $\mathbf{A}_{h}\in\mathbb{R}^{L}$ of each head are first broadcast along the feature dimension to form $\mathbf{A}\in\mathbb{R}^{L\times d}$, which is then multiplied with the value representations $\mathbf{V}_{h}$ (element-wise) to produce head-specific outputs, $\mathbf{O}_{h}\in\mathbb{R}^{L\times d}=\mathbf{A}\odot\mathbf{V}_{h}$. We concatenate these outputs across all heads to obtain $\mathbf{O}_{r}\in\mathbb{R}^{K\times L\times d}$. We then use a $2$-layer perceptron (MLP, $d$-dimensional layers) with Gaussian Error Linear Unit (GELU) activation to further capture nonlinearities. This is connected to a residual layer to form the output $\mathbf{O}_{m}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{r}+\texttt{MLP}(\mathbf{O}_{r})$. We then multiply $\mathbf{O}_{m}$ with a weight matrix $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times d}$ to obtain the output of MVCA, $\mathbf{O}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{m}\mathbf{W}_{\text{out}}$.

### v2

@@ Abstract
Multivariable time series forecasting methods can integrate information from exogenous variables, leading to significant prediction accuracy gains. The transformer architecture has been widely applied in various time series forecasting models due to its ability to capture long-range sequential dependencies. However, a naïve application of transformers often struggles to effectively model complex relationships among variables over time. To mitigate against this, we propose a novel architecture, termed Spectral Operator Neural Network (Sonnet). Sonnet applies learnable wavelet transformations to the input and incorporates spectral analysis using the Koopman operator. Its predictive skill relies on the Multivariable Coherence Attention (MVCA), an operation that leverages spectral coherence to model variable dependencies. Our empirical analysis shows that Sonnet yields the best performance on $34$ out of $47$ forecasting tasks with an average mean absolute error (MAE) reduction of $2.2\%$ against the most competitive baseline. We further show that MVCA can remedy the deficiencies of naïve attention in various deep learning models, reducing MAE by $10.7\%$ on average in the most challenging forecasting tasks.  [^1] Accepted for oral presentation at AAAI 2026. This is an extended version of the conference paper.
Code & datahttps://github.com/ClaudiaShu/Sonnet
@@ 1 Introduction
Multivariable time series (MTS) forecasting methods learn from multiple input variables to predict a single target variable (hidalgo2013multivariate). MTS models are deployed in many real-life applications, such as financial modelling (pniemira2004analytic; antunes2018forecasting; wtaylor2003using; fwitt1995forecasting), numerical weather prediction (fildes2011validation; cyoung2018data; dalton2022exogenous), and computational epidemiology (freyerdugas2013influenza; shaman2012forecasting; silva2020forecasting; morris2023neural). Recent machine learning models for time series forecasting tend to emphasise multivariate formulations (nie2023time; zeng2023are; liu2023koopa; ansari2024chronos; das2024decoder). While such models provide important insights for integrating traditional time series forecasting into deep learning techniques, e.g. capturing seasonality (lin2024cyclenet), frequency-domain methods (zhou2022fedformer; zhou2022film), and autoregression (zeng2023are; nie2023time), they often overlook the benefits of leveraging external information through exogenous indicators.
@@
In fact, empirical evidence suggests that capturing dependencies from external variables can increase the risk of overfitting for complex models (nie2023time). Moreover, existing methods that capture inter-variable dependencies (zhang2023crossformer; wang2024timemixer) have, in certain forecasting tasks, been outperformed by models that do not (lin2024cyclenet; luo2024deformabletst). However, this can be attributed to the choice of forecasting benchmarks, which, in many occasions, hold strong seasonal patterns (lin2024cyclenet; see also Appendix E.5), that can be effectively modelled using simpler autoregressive methods (zeng2023are; lin2024cyclenet). Additionally, models that do not capture inter-variable dependencies benefit from learning using longer look-back windows without significantly increasing modelling complexity and computational cost (han2023capacity), making them less likely to overfit. Nevertheless, these empirical outcomes are not necessarily valid for multivariable forecasting scenarios, especially for tasks where exogenous variables have strong predictive power. In such cases, explicitly modelling interactions across variables facilitates better performance (wang2024timexer; shu2025deformtime).
@@
Prior work has explored various ways in capturing inter-variable dependencies for multivariate or MTS forecasting tasks. Methods applying attention (vaswani2017attention) across variables (liu2024itransformer; ilbert2024samformer) can capture nonlinear dependencies, but disrupt temporal information by embedding sequences along the time dimension. Crossformer (zhang2023crossformer) attempts to rectify this by proposing a modified transformer structure where input is split into time series patches allowing to capture dependencies across both the time and variable dimensions at the sub-sequence level. ModernTCN (luo2024moderntcn), on the other hand, uses a convolutional kernel over both the variable and time dimension to jointly capture the inter- and intra-variable dependencies. However, both Crossformer and ModernTCN suffer from GPU computational overhead as the number of exogenous variables or the length of the input time series increases (shu2025deformtime; zhou2024efficanet). TimeXer (wang2024timexer) models dependencies between exogenous variables and the target variable separately, and then joins their learned embeddings with a cross-attention module. DeformTime (shu2025deformtime) yields superior results by using deformable attention to incorporate information from exogenous variables at different time steps. However, it only captures inter-variable dependencies within the reception field of a convolutional kernel, which limits the exploration of a wider range of exogenous variables.
@@
Frequency-based analysis techniques, including those using Fourier (bochner1953fourier; sorensen1987realvalued) or wavelet transform (varanini1997spectral; zhang2003detection; arneodo1988wavelet; farge1992wavelet; yu2024adawavenet), has been widely used in statistical approaches for identifying periodic patterns (priestley2018spectral) as well as in machine learning methods for forecasting (lange2021fourier; zhou2022fedformer; zhou2022film; piao2024fredformer). These methods compress temporal information (due to the Fourier transform) while some of them focus on capturing intra-variable dependencies (zhou2022fedformer; zhou2022film). Wavelet transform, albeit dependent on the chosen mother wavelet (demoortel2004wavelet; ngui2013wavelet), can maintain both time and frequency information. In the frequency domain, spectral coherence (white1990cross) serves as a powerful tool for capturing the correlation between variables at different frequencies (stein1972frequency; mima1999corticomuscular). To improve input projections into the frequency domain, lange2021fourier used a Koopman operator (mezic2005spectral; rowley2009spectral; avila2020datadriven; liu2023koopa), i.e. a spectral function that enables linear modelling of nonlinear changes to better address nonlinearities. More recently, AdaWaveNet (yu2024adawavenet) decomposed the input series into seasonal and long-term components, and used the wavelet transform to capture periodic information. This method accounts for inter-variable dependencies only over the seasonal components.
@@
Motivated by the aforementioned remarks, we propose Spectral Operator Neural Network (Sonnet), a model that captures MTS dependencies in the spectral domain using a learnable wavelet transform. Sonnet captures both intra- and inter-variable dependencies using a novel frequency-domain Multivariable Coherence Attention (MVCA) layer. Furthermore, it deploys a learnable Koopman operator for linearised transitions of temporal states (li2020learning). MVCA demonstrates stand-alone effectiveness when integrated into existing architectures, outperforming naïve attention and other modified attention mechanisms. Our key contributions are:
@@
    We assess forecasting accuracy on carefully curated MTS data sets, including established benchmarks complemented by tasks on weather forecasting, influenza prevalence, electricity consumption, and energy prices. The weather prediction and influenza prevalence tasks support a more thorough evaluation as they include a substantial amount of exogenous variables, multiple years ($\geq 10$) for training, and multiple test periods ($\geq 3$).
@@
    Sonnet reduces mean absolute error (MAE) by $2.2\%$ on average compared to the most performant baseline model (stat. sig., $p<10^{-3}$). In the more challenging tasks of influenza and weather modelling, MAE is reduced by $3.5\%$ and $2\%$, respectively. Performance gains persist as the forecasting horizon increases, demonstrating the effectiveness of Sonnet in longer-term forecasting.
@@ 2 MTS Forecasting Task Definition
We focus exclusively on multivariable time series (MTS) forecasting, whereby multiple input variables are used to predict a single target variable. We note that although for some baseline models multiple output variables may be present (multivariate forecasting), our evaluation is restricted to the prediction of one specific output. All models are trained and evaluated under a rolling window setup, with a look-back window and a forecasting horizon of $L$ and $H$ time steps, respectively. At each time step $t$, the $C$ observed exogenous variables over $L$ past time steps, $\{t-L+1,\dots,t\}$, are captured in an input matrix $\mathbf{X}_{t}\in\mathbb{R}^{L\times C}$. The autoregressive signal for the target (endogenous) variable is denoted by $\mathbf{y}_{t-\delta}\in\mathbb{R}^{L}$; this encompasses time steps $\{t-\delta-L+1,\dots,t-\delta\}$, where $\delta\in\mathbb{N}_{0}$ is an optional delay applied when the target variable is observed with a temporal lag. We capture the exogenous and endogenous input variables in $\mathbf{Z}_{t}=[\mathbf{X}_{t},\mathbf{y}_{t-\delta}]\in\mathbb{R}^{L\times(C+1)}$. The goal is to predict the target variable at time step $t+H$, where $H$ denotes the forecasting horizon. Hence, the output of a forecasting model is denoted by $\mathbf{y}_{t+H}\in\mathbb{R}^{H}$ and holds forecasts for time steps $\{t+1,\dots,t+H-1,t+H\}$. For models that conduct multivariate forecasting, the output includes predictions for all covariates and hence is denoted by $\mathbf{Y}_{t+H}=[\mathbf{X}_{t+H},\mathbf{y}_{t+H}]\in\mathbb{R}^{H\times(C+1)}$. The forecasting task is to learn $f:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Performance is measured based on the endogenous forecast at time step $t+H$, i.e. the last (temporally) element of $\mathbf{y}_{t+H}$, $y_{t+H}\in\mathbb{R}$. For notational simplicity, we omit temporal subscripts and use $\mathbf{Z}$ for $\mathbf{Z}_{t}$, $\mathbf{y}$ for $\mathbf{y}_{t-\delta}$, and $\mathbf{X}$ for $\mathbf{X}_{t}$.
@@ 3 Spectral Coherence with Sonnet
In this section, we provide a detailed description of our proposed model, Sonnet. It operates in the spectral domain via a learnable wavelet transform and introduces a Multivariable Coherence Attention (MVCA) module to capture both inter- and intra-variable dependencies. We further use a Koopman projection layer that enables stable temporal evolution via a learned linear operator.
3.1 Joint Embedding of Input Variables
@@ 3.2 Learnable Wavelet Transform
We then transform the time series embedding into the wavelet space that contains both time and frequency information, to capture both fine-grained details and overall trends across various temporal resolutions (mallat1989theory; daubechies1990wavelet). Specifically, after obtaining the input time series embedding $\mathbf{E}$, we first define a set of $K$ learnable wavelet transformations (also referred to as atoms), with the $k$-th atom held in a matrix $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ derived by
@@
where $\mathbf{t}\in\mathbb{R}^{L}$ is a row vector capturing normalised time steps with $\mathbf{t}_{i}=i/(L-1)$ for $i=0,\dots,L-1$, and $\mathbf{w}_{\alpha}$, $\mathbf{w}_{\beta}$, and $\mathbf{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable weight vectors that control the shape of the wavelet, each initialised randomly from a normal distribution. In particular, $\mathbf{w}_{\alpha}$ controls the width of the Gaussian envelope, and $\mathbf{w}_{\beta}$, $\mathbf{w}_{\gamma}$ respectively determine the linear or quadratic frequency modulation of the generated cosine waveforms. This formulation enables the atoms to adapt to localised time-frequency structures in the data. The time series embedding $\mathbf{E}\in\mathbb{R}^{L\times d}$ is then transformed into the wavelet space by projecting it onto each of the wavelet atoms $\mathbf{M}_{k}\in\mathbb{R}^{d\times L}$ using $\mathbf{P}_{k}=\mathbf{E}\odot\mathbf{M}_{k}^{\top}$, where $\mathbf{P}_{k}\in\mathbb{R}^{L\times d}$ denotes the embedding’s projection for the $k$-th atom, and $\odot$ is element-wise multiplication. The transformed wavelet across all $K$ atoms is denoted by $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$. The aforementioned steps help to preserve temporal structure while decomposing the input into multi-resolution time-frequency components using adaptive wavelet transforms that can capture both short- and long-term patterns in the data.
@@ 3.3 Multivariable Coherence Attention (MVCA)
Given the obtained input embedding matrix in wavelet space $\mathbf{P}\in\mathbb{R}^{K\times L\times d}$, where $L$ and $d$ are the time and variable dimension, $K$ is the number of wavelets after transformation, we first linearly project it to query, key and value embeddings, denoted as $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}\in\mathbb{R}^{K\times L\times d}$, using weight matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ respectively, as in $\mathbf{Q}=\mathbf{P}\mathbf{W}_{q}$. We then consider the embeddings from different wavelet transformations (indexed by $k$) as separate attention heads. Specifically, for each attention head, we obtain the sub-tensor of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, denoted as $\mathbf{Q}_{h}$, $\mathbf{K}_{h}$, and $\mathbf{V}_{h}\in\mathbb{R}^{L\times d}$, as the query, key, and value embeddings. Therefore, each head captures a distinct subspace of the original embedding, where subspaces correspond to different wavelet transformations of the input. Including a multi-head structure enables the model to learn diverse dependencies in parallel (vaswani2017attention).
We then apply a Fast Fourier Transform (FFT) along the variable dimension (dudgeon1984multidimensional) of each transformer head to transfer the query and key embeddings to the frequency domain, i.e. $\mathbf{Q}_{f}=\mathrm{FFT}\left(\mathbf{Q}_{h}\right)$ and $\mathbf{K}_{f}=\mathrm{FFT}\left(\mathbf{K}_{h}\right)$, with both $\mathbf{Q}_{f}$ and $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$, where $\ell=\lfloor\frac{d}{2}\rfloor+1$. Each frequency bin in the transformed data ($\mathbf{Q}_{f}$, $\mathbf{K}_{f}$) contains information from all original inputs to the FFT (lee-thorp2022fnet), capturing both fine-grained (high-frequency) and global (low-frequency) patterns. The cross-spectral density $\mathbf{P}_{qk}\,\in\,\mathbb{C}^{L\times\ell}$ and power-spectral densities $\mathbf{P}_{qq}$, $\mathbf{P}_{kk}\,\in\,\mathbb{R}^{L\times\ell}$ are obtained as follows:
$\mathbf{P}_{qk}=\mathbf{Q}_{f}\odot\mathbf{K}_{f}^{\ast}\,,\,\,\,\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}\,,\,\,\,\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}\,$, (2)
@@
where $\epsilon=10^{-6}$ mitigates division by 0. $\mathbf{C}_{qk}$ captures the linear dependency between sequences across frequency bands. Therefore, the coherence here assigns higher importance to the time steps where the query and key hold more similar averaged values across multiple frequencies.
Following the common design of attention layers (vaswani2017attention), we scale, normalise, and regularise $\mathbf{C}_{qk}$, i.e. $\mathbf{A}_{h}=\mathrm{Dropout}(\mathrm{Softmax}(\mathbf{C}_{qk}/\sqrt{d}))$. The attention weights $\mathbf{A}_{h}\in\mathbb{R}^{L}$ of each head are first broadcast along the feature dimension to form $\mathbf{A}\in\mathbb{R}^{L\times d}$, which is then multiplied with the value representations $\mathbf{V}_{h}$ (element-wise) to produce head-specific outputs, $\mathbf{O}_{h}\in\mathbb{R}^{L\times d}=\mathbf{A}\odot\mathbf{V}_{h}$. We concatenate these outputs across all heads to obtain $\mathbf{O}_{r}\in\mathbb{R}^{K\times L\times d}$. We then use a $2$-layer perceptron ($\mathrm{MLP}$, $d$-dimensional layers) with Gaussian Error Linear Unit (GELU) activation to further capture nonlinearities. This is connected to a residual layer to form the output $\mathbf{O}_{m}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{r}+\mathrm{MLP}(\mathbf{O}_{r})$. We then multiply $\mathbf{O}_{m}$ with a weight matrix $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times d}$ to obtain the output of MVCA, $\mathbf{O}\in\mathbb{R}^{K\times L\times d}=\mathbf{O}_{m}\mathbf{W}_{\text{out}}$.